Trade Kings, one of the biggest companies in Zambia, has hired you to advise them on pricing policy. One of the things the company would like to know is how much a 1.5 percent increase in prices of all the goods they are selling arelikely to affect sales. What critical information you have to have in order to appropriately advice the company? Explain why these facts are important.



1
Expert's answer
2022-03-01T14:50:40-0500

What critical information do you have to have in order to appropriately advise the company? Explain why these facts are important.

You need to know the price elasticity of demand before setting the price. Elastic demand is a term used to describe a situation in which demand changes in response to a small change in price. The percentage change in quantity divided by the percentage change in price yields the price elasticity of demand. Therefore, apart from the percentage change in the price of the good, you need to calculate the percentage change in the quantity.

The formula of the price elastic of demand is given by:

PED = abs["\frac{\%\,Δ\ Quantity\ demanded}{\%\,Δ\ price}" ]

where;

% Δ Qd = "\frac{Q_1- Q_0 }{Q_0}"


% Δ P = "\frac{P_1- P_0 }{P_0}"


  • Q1 is the new quantity while Q0 is the initial quantity.
  • P1 is the new price while P0 is the initial price.

Interpretation:

  • If −(elasticity of demand) > 1, demand is relatively elastic.
  • If −(elasticity of demand) < 1, demand is relatively inelastic.
  • It is an inferior good if the income elasticity of demand is negative.
  • It is a normal good if the demand elasticity of income is positive.
  • It is a luxury good if the demand elasticity of income is larger than one.